OpenStudy (anonymous):

f(g(-2)) f(x)=-x^3 g(x)= 1/2x^2 can anyone give me the answer and help me..

OpenStudy (jiteshmeghwal9):

g(x).=1/2x^2 by putting the value of 'x' in function g(-2)=1/2(-2)^2=?

OpenStudy (kc_kennylau):

\[\large\begin{array}{cl} &f(g(2)\\ =&f\left[\frac12(2^2)\right]\\ =&f(2)\\ =&-2^3\\ =&-8 \end{array}\]

OpenStudy (kc_kennylau):

P.S. \(g(2)=\frac122^2=2\)

OpenStudy (anonymous):

thank youso much :) could you help me with.. f(x) + g(x) f(x)=3x+4 g(x)=7x-9

OpenStudy (kc_kennylau):

Just freaking add them together :)

OpenStudy (anonymous):

hahah oh my gosh,, im dumb..